OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by y= 1/x^5, y=0, x=2, x=9 about the -axis. please help with the answer i have tried and i keep gettign wrong answer :(

OpenStudy (alekos):

whats your integral?

OpenStudy (anonymous):

When x=2, y=1/32. When x=9, y approaches zero. Approximately, \[A=\int\limits_{0}^{1/32}π[f ^{-1}(y)]^2dy\]\[=π \int\limits_{0}^{1/32}[y ^{-1/5}]^2dy\]\[=π[y ^{3/5}/(3/5)] from 0 1/32\]\[=\frac{ 3π}{ 5 }[0-(\frac{ 1 }{ 32 })^{3/5}]\] Since we want the magnitude, \[A = \frac{ 3π }{ 5\times32^{3/5} }\] = 0.236 Is it right?